# Impact of Valley Degeneracy on Thermoelectric Properties of Zigzag Graphene Nanoribbons with Staggered Sublattice Potentials and Transverse Electric Fields

### v1

@@ I Introduction
The landmark discovery of graphene in 2004 by Novoselov and Geim [Novoselovks, 24] prompted investigations into quasi-one-dimensional electronics, optoelectronics, and thermoelectric devices utilizing nanowires formed from two-dimensional materials [WangHM, 25,Chunhua, 26]. Graphene nanoribbons (GNRs) have garnered attention due to their low thermal conductance, as reported in theoretical calculations [OuyangYJ, 15], owing to their atomic-scale thickness. Edge-disordered zigzag GNRs (ZGNRs) [SevincliH, 16], semiconducting armchair GNRs (AGNRs) [MazzamutoF, 17], and GNRs with topological states [XuY, 22] exhibit significantly enhanced $ZT$ values compared to gapless graphenes [ChangPH, 23]. These GNRs achieve $ZT$ values greater than one, although their power factors are relatively small. TE devices require $ZT$ values exceeding three, along with optimized electrical power outputs [Whitney, 27]. Enhancing the power factor values through increased level degeneracy has been a focal point of numerous studies [Kuo1, 28Ranganayakulu, 36]. Materials like $GeTe$ exhibit very high $ZT$ at elevated temperatures, where valley degeneracy may play a crucial role [Toriyama, 32Ranganayakulu, 36].
@@
While bottom-up techniques provide precise atomic control over various GNRs, the existence of semiconducting GNRs with valley degeneracy has not yet been reported [Cai, 37JeilJ, 55]. To investigate the impact of valley degeneracy on the thermoelectric properties of GNRs, we propose utilizing ZGNRs with staggered sublattice potentials. These potentials can be implemented by coupling the ZGNRs to substrates such as silicon carbide or hexagonal boron nitride [ZhaoJ, 56]. We consider three configurations of ZGNRs with different staggered sublattice potentials denoted as (a) w-AB/n-C/w-AB, (b) w-AB/n-C/w-BA, and (c) w-BA/n-C/w-AB, where integers w and n represent the widths of ZGNRs with and without sublattice potential, respectively. The w-AB/4-C/w-AB structures exhibit semiconducting behavior and valley degeneracy in two valleys within the conduction and valence subbands. This valley degeneracy arises from band inversion of the flat bands induced by inversion symmetry breaking in the transverse direction, facilitated by the underlying substrates. The primary objective of this study is to investigate how valley degeneracy affects the thermoelectric properties of ZGNRs supported on these substrates.
@@
Figure 1: Schematic diagram illustrating three types of ZGNRs with staggered sublattice potentials: (a) 4-AB/4-C/4-AB, (b) 4-AB/4-C/4-BA, and (c) 4-BA/4-C/4-AB. In (a), ZGNR without inversion symmetry in the y direction. The square bracket indicates the unit cell of 4-AB/4-C/4-AB. In (b) and (c), ZGNRs remain the inversion symmetry in the y direction. Atoms marked with blue and white colors indicate these carbon atoms on the boron nitride substrates. Diagram (d) depicts the armchair edges of the 4-AB/4-C/4-AB segment in contact with electrodes. Symbols $\Gamma_{L}$ ($\Gamma_{R}$) represent the electron tunneling rate between the left (right) electrode and the leftmost (rightmost) atoms at the armchair edges, and $T_{L}$ ($T_{R}$) denotes the equilibrium temperature of the left (right) electrode. $F_{y}$ denotes a uniform transverse electric field along the armchair direction.
@@ II Calculation Methodology
Here, $E_{\ell,j}$ represents the on-site energy of the orbital in the ${\ell}$-th row and $j$-th column. The operators $d^{\dagger}_{\ell,j}$ and $d_{\ell,j}$ create and annihilate an electron at the atom site denoted by ($\ell$,$j$). The parameter $t_{(\ell,j),(\ell^{\prime},j^{\prime})}$ characterizes the electron hopping energy from site ($\ell^{\prime}$,$j^{\prime}$) to site ($\ell$,$j$). We set $t_{(\ell,j),(\ell^{\prime},j^{\prime})}=t_{pp\pi}=2.7$ eV for the nearest-neighbor hopping strength. Because we have considered ZGNRs on the hexagonal boron nitride substrates, staggered sublattice potentials are $E_{A}=\delta$ and $E_{B}=-\delta$, which are relative to $p_{z}$ orbital of carbon atom $E_{C}=0$[QiaoZH, 57]. Note that if one stack ZGNRs on other two dimensional material substrates and then spin orbital coupling may not be ignored[LuCP, 58].
In the linear response region, the electrical conductance ($G_{e}$), Seebeck coefficient ($S$) and electron thermal conductance ($\kappa_{e}$) can be computed using $G_{e}=e^{2}{\cal L}_{0}$, $S=-{\cal L}_{1}/(eT{\cal L}_{0})$ and $\kappa_{e}=\frac{1}{T}({\cal L}_{2}-\frac{{\cal L}^{2}_{1}}{{\cal L}_{0}})$ with ${\cal L}_{n}$ ($n=0,1,2$) defined as
@@
Here, $f(\varepsilon)=1/(1+\exp((\varepsilon-\mu)/k_{B}T))$ represents the Fermi distribution function of electrodes at equilibrium chemical potential $\mu$. The constants $e$, $h$, $k_{B}$, and $T$ denote the electron charge, Planck’s constant, Boltzmann’s constant, and the equilibrium temperature of the electrodes, respectively. ${\cal T}_{LR}(\varepsilon)$ signifies the transmission coefficient of a ZGNR connected to electrodes, and it can be calculated using the formula ${\cal T}_{LR}(\varepsilon)=4Tr[\Gamma_{L}(\varepsilon)G^{r}(\varepsilon)\Gamma_{R}(\varepsilon)G^{a}(\varepsilon)]$ [Kuo3, 30,SunQF, 59], where $\Gamma_{L}(\varepsilon)$ and $\Gamma_{R}(\varepsilon)$ denote the tunneling rate (in energy units) at the left and right leads, respectively, and ${G}^{r}(\varepsilon)$ and ${G}^{a}(\varepsilon)$ are the retarded and advanced Green’s functions of the GNRs, respectively. The tunneling rates ($\Gamma_{L(R)}(\varepsilon)=Im\sum^{r}_{L(R)}(\varepsilon)$) are determined by the imaginary part of the self-energy originating from the coupling between the left (right) electrode and its adjacent GNR atoms. In terms of tight-binding orbitals, $\Gamma_{\alpha}(\varepsilon)$ and Green’s functions are matrices. For simplicity, $\Gamma_{\alpha}(\varepsilon)$ for interface atoms possesses diagonal entries with a common value of $\Gamma_{t}$ [Kuo3, 30]. Contact properties between 2D nanostructures and metallic electrodes typically induce Schottky barriers [Areshkin, 60Matsuda, 64], posing challenges in achieving tunable Fermi energy due to Fermi energy pinning [ChenRS, 65]. Despite numerous theoretical studies striving to elucidate this crucial behavior from first principles, the theoretical limitations result in obtaining only qualitative results regarding $\Gamma_{t}$ arising from the contact junction [LeeG, 52].
@@ III.1 Electronic Structures of ZGNRs with staggered sublattice potentials
To depict the electronic properties of ZGNRs with staggered sublattice potentials, as shown in Fig. 1(a)-1(c), we calculated the electronic structures of three configurations: 4-AB/4-C/4-AB, 4-BA/4-C/4-AB, and 4-AB/4-C/4-BA, based on the Hamiltonian in Eq. (1). These electronic structures are presented in Fig. 2(a), 2(b), and 2(c), respectively. Among ZGNRs with sublattice potentials, only the w-AB/4-C/w-AB structure exhibits a semiconducting phase. Notably, the w-AB/4-C/w-AB structure also displays valley degeneracy in the frontier conduction and valence subbands, albeit with shallow valley depths. In contrast, both the 4-BA/4-C/4-AB and 4-AB/4-C/4-BA configurations exhibit metallic phases, characterized by asymmetrical electronic structures distinct from those of ZGNRs without staggered sublattice potentials, where symmetrical flat bands are observed near the charge neutral point (CNP) [Nakada, 66,Wakabayashi, 67].
Figures 2(d), 2(e), and 2(f) illustrate the electronic structures of w-AB/4-C/w-AB for three different widths, where w $=8,12$ and $16$. The valley structure of the frontier conduction and valence subbands is sensitive to the widths of ZGNRs coupled to the substrates. Based on the findings of Fig. 2, a semiconducting phase with valley degeneracy emerges from the w-AB/4-C/w-AB structure due to broken inversion symmetry in the transverse direction. However, this valley structure is suppressed by the coupling between the frontier and secondary conduction (valence) subbands. The mechanism of valley degeneracy in topological thermoelectric materials remains a topic of current interest [ToriyamaMY, 68-ZhuZY, 70].
@@
When transverse electric fields $F_{y}$ are applied to ZGNRs, the electric potential $U=eF_{y}y$ affects each site of the ZGNRs [ZhaoFZ, 71]. Here, $F_{y}=V_{y}/L_{a}$, where $V_{y}$ is the applied bias and $L_{a}$ is the width of the ZGNRs. We investigate the electronic structures of the 4-AB/4-C/4-AB structure under transverse electric fields, as depicted in Fig. 3. Increasing $V_{y}$ results in a reduction of the band gap in the 4-AB/4-C/4-AB structure. At $V_{y}=0.9$ V , the electronic structures of the frontier conduction and valence subbands exhibit flat bands. Conversely, negative values of $V_{y}$ reveal more pronounced valley structures in Fig. 3(d), 3(e), and 3(f). This negative $V_{y}$ enhances the inversion symmetry breaking of ZGNRs [ZhuZY, 70]. The directionality of the electric field influences the valley structure of 4-AB/4-C/4-AB ZGNRs with asymmetrical staggered sublattice potentials. Such an electric field-dependent valley structure could potentially find applications in valleytronics [Schaibley, 72-Motohiko, 75].
@@ III.2 Transmission Coefficient of 4-AB/4-C/4-AB under Transverse Electric Fields
Synthesizing ZGNRs longer than 20 nm using bottom-up techniques remains challenging [LlinasJP, 76, JacobsePH, 77]. Therefore, investigating the size effect of ZGNR segments on valley degeneracy is crucial. Here, we present the energy levels of ZGNR segments with $N_{z}=81$ ( $L_{z}=9.84$nm) as functions of the sublattice potential magnitude $\delta$ under various transverse electric fields in Fig. 4. When $\delta=0$ and $V_{y}=0$, metallic ZGNRs exhibit discrete energy levels due to their finite sizes, as seen in Fig. 4(a). The energy levels $E_{HO}$ (highest occupied molecular orbital, HOMO) and $E_{LU}$ (lowest unoccupied molecular orbital, LUMO) are very close to the charge neutral point ($E=0$) due to small overlap between the wave functions of zigzag edge states. As $\delta$ increases, $E_{HO}$ and $E_{LU}$ separate. Moreover, the number of level degeneracies, indicated by a wider range of levels with stronger color intensity, increases for $\delta>0.45$ eV. These results in Fig. 4(a) can be understood in the context of the electronic structures of 4-AB/4-C/4-AB ZGNRs shown in Appendix A (Fig. A.1). The extent of level degeneracy in 4-AB/4-C/4-AB ZGNR segments corresponds to the depth of valley structures in infinite 4-AB/4-C/4-AB ZGNRs. Upon applying transverse electric fields, the patterns of energy levels with respect to $\delta$ change dramatically, as illustrated in Figs. 4(b), 4(c), and 4(d). The range of energy level degeneracy with respect to transverse electric fields aligns with the electronic structures shown in Fig. 3. Notably, while ZGNRs without staggered sublattice potential have demonstrated valley structures under strong transverse electric fields, the extent of level degeneracy is limited, and the electric field-direction-dependent valley structure disappears [LeeKW, 78]
@@
Since the transmission coefficient spectra of nanostructures are influenced by the wave-functions of their energy levels [Kuo3, 30], understanding the wave functions of each energy level for 4-AB/4-C/4-AB ZGNR segments with various $\delta$ values is crucial. In Fig. 5(a), the probability densities $|\Psi(\ell,j)(E=0)|^{2}$ demonstrate localized and symmetrical zigzag edge states in the $y$ direction for $\delta=0$ eV and $V_{y}=0$. Figures 5(b) and 5(c) show the probability densities of $E_{HO}=-0.253$ eV and $E_{LU}=0.253$ eV, respectively, depicting the wave functions of the HOMO and the LUMO at $\delta=0.27$ eV. Analysis of their probability density distributions reveals that $E_{HO(LU)}=\mp 0.253$ eV are primarily determined by $E_{B}$ and $E_{A}$, respectively. The maximum probability densities $|\Psi(\ell,j)(E_{HO})|^{2}$ and $|\Psi(\ell,j)(E_{LU})|^{2}$ are concentrated at $j=12$ and $j=1$, respectively, with $|\Psi(\ell,j)(E_{HO})|^{2}$ being vanishingly small for $j=7,9$ and $11$, and $|\Psi(\ell,j)(E_{LU})|^{2}$ for $j=2,4$ and $6$. In Fig. 5(d), 5(e) and 5(f) at $\delta=0.9$ eV, we present the probability density distributions of $E_{1}=-0.704$ eV, $E_{2}=-0.7053$ eV and $E_{3}=-0.7284$ eV, which are below the CNP. These states exhibit delocalized and asymmetrical characteristics in the $y$ direction, indicative of band inversion in flat bands for ZGNRs. The probability densities $|\Psi(\ell=1,j)(E_{i})|^{2}$ and $|\Psi(\ell=81,j)(E_{i})|^{2}$ significantly influence the transmission coefficient spectra of 4-AB/4-C/4-AB ZGNR segments with armchair edges coupled to the electrodes shown in Fig. 1(d).
@@
According to Eq. (2), the thermoelectric quantities of ZGNRs, including electrical conductance, Seebeck coefficient, and electron thermal conductance, are determined by the area and shape of the transmission coefficient (${\cal T}_{LR}(\varepsilon)$) curve. Therefore, it is necessary to clarify the effects of contact properties on ${\cal T}_{LR}(\varepsilon)$. In Fig. 6, we present the calculated ${\cal T}_{LR}(\varepsilon)$ curves of 4-AB/4-C/4-AB ZGNR segments for various $\Gamma_{t}$ values, which depend on the contact geometries and the electrode materials. For simplicity of analysis, $\varepsilon\leq 0$ is considered throughout this article. As seen in Fig. 6(a), the ${\cal T}_{LR}(\varepsilon)$ curve shows many discrete peaks under a weak coupling strength of $\Gamma_{t}=0.27$ eV. With increasing $\Gamma_{t}$ values, the area of ${\cal T}_{LR}(\varepsilon)$ curve increases, and the shapes of these areas also change. For instance, in Fig. 6(b), the area of ${\cal T}_{LR}(\varepsilon)$ curve shows a right-triangle shape with a steep change with respect to $\varepsilon$ on the side toward the central gap. In Figs. 6(c) and 6(d), the areas of ${\cal T}_{LR}(\varepsilon)$ curves show rectangular shapes. For $\Gamma_{t}=2.7$ eV, we observe an arch shape in Fig. 6(f). The level degeneracy is characterized by ${\cal T}_{LR}(\varepsilon)$ with values near two. For $\Gamma_{t}=2.7$ eV, the electrode materials can be regarded as graphenes since $\Gamma_{t}=t_{pp\pi}$.
@@
Figure 6: Transmission coefficient ${\cal T}_{LR}(\varepsilon)$ of 4-AB/4-C/4-AB ZGNR segments as functions of $\varepsilon$ for various $\Gamma_{t}$ values at $N_{a}=12$, $N_{z}=81$, and $V_{y}=-0.9$ V: (a) $\Gamma_{t}=0.27$ eV, (b) $\Gamma_{t}=0.45$ eV, (c) $\Gamma_{t}=0.63$ eV, (d) $\Gamma_{t}=0.81$ eV, (e) $\Gamma_{t}=0.99$ eV, and (f) $\Gamma_{t}=2.7$ eV. Here, $N_{z}=81$ corresponds to a channel length of $L_{z}=9.84$ nm.

### v3

@@ I Introduction
The landmark discovery of graphene in 2004 by Novoselov and Geim [Novoselovks, 24] prompted investigations into quasi-one-dimensional electronics, optoelectronics, and thermoelectric devices utilizing nanowires formed from two-dimensional materials [WangHM, 25, Chunhua, 26]. Graphene nanoribbon (GNR)-based devices include field-effect transistors [PassiV, 27], spintronics [CaoL, 28], gas sensors [LiuGG, 29], nanoscale rotators [LiXB, 30], and electrical diodes [DingWC, 31]. GNRs have also garnered attention in thermoelectric applications due to their low thermal conductance, as reported in theoretical calculations [OuyangYJ, 15], which is attributed to their atomic-scale cross-section. Edge-disordered zigzag GNRs (ZGNRs) [SevincliH, 16], semiconducting armchair GNRs (AGNRs) [MazzamutoF, 17], and GNRs with topological states [XuY, 22] exhibit significantly enhanced $ZT$ values compared to gapless graphenes [ChangPH, 23]. These GNRs achieve $ZT$ values greater than one, although their power factors are relatively small. TE devices require $ZT$ values exceeding three, along with optimized electrical power outputs [Whitney, 32]. Enhancing the power factor values through increased level degeneracy has been a focal point of numerous studies [Kuo1, 33Ranganayakulu, 41]. Materials like $GeTe$ exhibit very high $ZT$ at elevated temperatures, where valley degeneracy may play a crucial role [Toriyama, 37Ranganayakulu, 41].
@@
While bottom-up techniques provide precise atomic control over various GNRs, the existence of semiconducting GNRs with valley degeneracy has not yet been reported [Cai, 42JeilJ, 60]. To investigate the impact of valley degeneracy on the thermoelectric properties of GNRs, we propose utilizing ZGNRs with staggered sublattice potentials. These potentials can be implemented by coupling the ZGNRs to substrates such as silicon carbide or hexagonal boron-nitride [ZhaoJ, 61-ZhouSY, 63].Most recently, an experimental report indicated a band gap of $E_{g}=0.6$ eV for graphene, attributed to the effect of silicon carbide substrates [ZhaoJ, 61]. This impressive $E_{g}$ value is significantly larger than theoretical predictions [Giovannetti, 62, ZhouSY, 63]. In Fig. 1, we consider three scenarios of zigzag graphene nanoribbons (ZGNRs) with staggered sublattice potentials, denoted as (a) w-AB/n-C/w-AB, (b) w-AB/n-C/w-BA, and (c) w-BA/n-C/w-AB. Here, the integers w and n represent the widths of ZGNRs with and without sublattice potential, respectively. Sublattices A and B are subjected to the staggered potentials. One of the crucial findings is that the w-AB/4-C/w-AB structures exhibit semiconducting behavior and valley degeneracy in two valleys within the conduction and valence subbands. This valley degeneracy arises from band inversion of the flat bands induced by inversion symmetry breaking in the transverse direction, facilitated by the underlying substrates. The primary objective of this study is to investigate how valley degeneracy affects the thermoelectric properties of ZGNRs supported on these substrates.
@@
Figure 1: Schematic diagram illustrating three types of ZGNRs with staggered sublattice potentials: (a) 4-AB/4-C/4-AB, (b) 4-AB/4-C/4-BA, and (c) 4-BA/4-C/4-AB. In (a), ZGNR without inversion symmetry in the y direction. The square bracket indicates the unit cell of 4-AB/4-C/4-AB. In (b) and (c), ZGNRs remain the inversion symmetry in the y direction. Atoms marked in blue and white indicate that these carbon atoms are subjected to staggered potentials $E_{A}=\delta$ and $E_{B}=-\delta$. Additionally, $E_{C}=0$ denotes the $p_{z}$ orbital of the carbon atom. Diagram (d) depicts the armchair edges of the 4-AB/4-C/4-AB segment in contact with electrodes. Symbols $\Gamma_{L}$ ($\Gamma_{R}$) represent the electron tunneling rate between the left (right) electrode and the leftmost (rightmost) atoms at the armchair edges, and $T_{L}$ ($T_{R}$) denotes the equilibrium temperature of the left (right) electrode. $F_{y}$ denotes a uniform transverse electric field along the armchair direction.
@@ II Calculation Methodology
Here, $E_{\ell,j}$ represents the on-site energy of the orbital in the ${\ell}$-th row and $j$-th column. The operators $d^{\dagger}_{\ell,j}$ and $d_{\ell,j}$ create and annihilate an electron at the atom site denoted by ($\ell$,$j$). The parameter $t_{(\ell,j),(\ell^{\prime},j^{\prime})}$ characterizes the electron hopping energy from site ($\ell^{\prime}$,$j^{\prime}$) to site ($\ell$,$j$). We set $t_{(\ell,j),(\ell^{\prime},j^{\prime})}=t_{pp\pi}=2.7$ eV for the nearest-neighbor hopping strength. Because we have considered ZGNRs on the hexagonal boron nitride substrates, staggered sublattice potentials are $E_{A}=\delta$ and $E_{B}=-\delta$, which are relative to $p_{z}$ orbital of carbon atom $E_{C}=0$[Giovannetti, 62,QiaoZH, 64]. Note that if one stack ZGNRs on other two dimensional material substrates and then spin orbital coupling may not be ignored[LuCP, 65]. When a uniform transverse electric field $E_{F}$ is introduced into the system, electrical potential $U=eF_{y}y$ affects each site of the ZGNRs. Here, $F_{y}=V_{y}/L_{a}$, where $V_{y}$ is the applied bias and $L_{a}$ is the width of the ZGNRs [Pizzochero, 66].
@@
[…]
Here, $f(\varepsilon)=1/(1+\exp((\varepsilon-\mu)/k_{B}T))$ represents the Fermi distribution function of electrodes at equilibrium chemical potential $\mu$. The constants $e$, $h$, $k_{B}$, and $T$ denote the electron charge, Planck’s constant, Boltzmann’s constant, and the equilibrium temperature of the electrodes, respectively. ${\cal T}_{LR}(\varepsilon)$ signifies the transmission coefficient of a ZGNR connected to electrodes, and it can be calculated using the formula ${\cal T}_{LR}(\varepsilon)=4Tr[\Gamma_{L}(\varepsilon)G^{r}(\varepsilon)\Gamma_{R}(\varepsilon)G^{a}(\varepsilon)]$ [Kuo3, 35,SunQF, 67], where $\Gamma_{L}(\varepsilon)$ and $\Gamma_{R}(\varepsilon)$ denote the tunneling rate (in energy units) at the left and right leads, respectively, and ${G}^{r}(\varepsilon)$ and ${G}^{a}(\varepsilon)$ are the retarded and advanced Green’s functions of the GNRs, respectively. The tunneling rates ($\Gamma_{L(R)}(\varepsilon)=Im\sum^{r}_{L(R)}(\varepsilon)$) are determined by the imaginary part of the self-energy originating from the coupling between the left (right) electrode and its adjacent GNR atoms. In terms of tight-binding orbitals, $\Gamma_{\alpha}(\varepsilon)$ and Green’s functions are matrices. For simplicity, $\Gamma_{\alpha}(\varepsilon)$ for interface atoms possesses diagonal entries with a common value of $\Gamma_{t}$ [Kuo3, 35]. Contact properties between 2D nanostructures and metallic electrodes typically induce Schottky barriers [Areshkin, 68Matsuda, 72], posing challenges in achieving tunable Fermi energy due to Fermi energy pinning [ChenRS, 73]. Despite numerous theoretical studies aiming to elucidate this crucial behavior from first principles, the inherent limitations result in only qualitative insights regarding $\Gamma_{t}$ arising from the contact junction [LeeG, 57,Matsuda, 72]. Therefore, the tight-binding model serves as an effective approach for calculating the electronic structures and charge transport of GNRs [Kuo1, 33-Kuo3, 35,JeilJ, 60,Pizzochero, 66]. In this work, we developed Fortran computational code to evaluate the electronic structures, transmission coefficients, and thermoelectric quantities of ZGNRs with staggered sublattice potentials and transverse electric fields.
@@ III.1 Electronic Structures of ZGNRs with Staggered Sublattice Potentials
To depict the electronic properties of ZGNRs with staggered sublattice potentials, as shown in Fig. 1(a)-1(c), we calculated the electronic structures of three configurations: 4-AB/4-C/4-AB, 4-BA/4-C/4-AB, and 4-AB/4-C/4-BA, based on the Hamiltonian in Eq. (1). These electronic structures are presented in Fig. 2(a), 2(b), and 2(c), respectively. Among ZGNRs with sublattice potentials, only the w-AB/4-C/w-AB structure exhibits a semiconducting phase. Notably, the w-AB/4-C/w-AB structure also displays valley degeneracy in the frontier conduction and valence subbands, albeit with shallow valley depths. In contrast, both the 4-BA/4-C/4-AB and 4-AB/4-C/4-BA configurations exhibit metallic phases, characterized by asymmetrical electronic structures distinct from those of ZGNRs without staggered sublattice potentials, where symmetrical flat bands are observed near the charge neutral point (CNP) [Nakada, 74,Wakabayashi, 75].
Figures 2(d), 2(e), and 2(f) illustrate the electronic structures of w-AB/4-C/w-AB for three different widths, where w $=8,12$ and $16$. The valley structure of the frontier conduction and valence subbands is sensitive to the widths of ZGNRs coupled to the substrates. Based on the findings of Fig. 2, a semiconducting phase with valley degeneracy emerges from the w-AB/4-C/w-AB structure due to broken inversion symmetry in the transverse direction. However, this valley structure is suppressed by the coupling between the frontier and secondary conduction (valence) subbands. The mechanism of valley degeneracy in topological thermoelectric materials remains a topic of current interest [ToriyamaMY, 76-ZhuZY, 78].
@@
When transverse electric fields $F_{y}$ are applied to ZGNRs, the electric potential $U=eV_{y}y/L_{a}$ affects each site of the ZGNRs [Pizzochero, 66]. This indicates that $E_{A}$, $E_{B}$ and $E_{C}$ depend on the strength and direction of $V_{y}$. We investigate the electronic structures of the 4-AB/4-C/4-AB structure under transverse electric fields, as depicted in Fig. 3. Increasing $V_{y}$ results in a reduction of the band gap in the 4-AB/4-C/4-AB structure. At $V_{y}=0.9$ V , the electronic structures of the frontier conduction and valence subbands exhibit flat bands. Conversely, negative values of $V_{y}$ reveal more pronounced valley structures in Fig. 3(d), 3(e), and 3(f). This negative $V_{y}$ enhances the inversion symmetry breaking of ZGNRs [ZhuZY, 78]. To illustrate reduction and enhancement of the band gap, let’s consider the sublattice sites at $j=1$ with energy level $E_{A}$ and $j=N_{a}$ with energy level $E_{B}$. For $V_{y}>0$, the sublattice potentials are modified to $\delta_{A}=\delta-eV_{y}/2$ and $\delta_{B}=-\delta+eV_{y}/2=-\delta_{A}$. For $V_{y}<0$, the potentials $E_{A}=\delta$ and $E_{B}=-\delta$ are adjusted to $\delta_{A}=\delta+eV_{y}/2$ and $\delta_{B}=-\delta-eV_{y}/2=-\delta_{A}$. The enhancement and reduction of the sublattice potentials directly influence the corresponding changes in the band gaps. The directionality of the electric field influences the valley structure of 4-AB/4-C/4-AB ZGNRs with asymmetrical staggered sublattice potentials. Such an electric field-dependent valley structure could potentially find applications in valleytronics [Schaibley, 80-Motohiko, 83].
@@ III.2 Transmission Coefficient of 4-AB/4-C/4-AB under Transverse Electric Fields
Synthesizing ZGNRs longer than 20 nm using bottom-up techniques remains challenging [LlinasJP, 84, JacobsePH, 85]. Therefore, investigating the size effect of ZGNR segments on valley degeneracy is crucial. Here, we present the energy levels of ZGNR segments with $N_{z}=81$ ( $L_{z}=9.84$nm) as functions of the sublattice potential magnitude $\delta$ under various transverse electric fields in Fig. 4. When $\delta=0$ and $V_{y}=0$, metallic ZGNRs exhibit discrete energy levels due to their finite sizes, as seen in Fig. 4(a). The energy levels $E_{HO}$ (highest occupied molecular orbital, HOMO) and $E_{LU}$ (lowest unoccupied molecular orbital, LUMO) are very close to the charge neutral point (CNP) due to small overlap between the wave functions of zigzag edge states. As $\delta$ increases, $E_{HO}$ and $E_{LU}$ separate. Moreover, the number of level degeneracies, indicated by a wider range of levels with stronger color intensity, increases for $\delta>0.45$ eV. These results in Fig. 4(a) can be understood in the context of the electronic structures of 4-AB/4-C/4-AB ZGNRs shown in Appendix A (Fig. A.1). The extent of level degeneracy in 4-AB/4-C/4-AB ZGNR segments corresponds to the depth of valley structures in infinite 4-AB/4-C/4-AB ZGNRs. Upon applying transverse electric fields, the patterns of energy levels with respect to $\delta$ change dramatically, as illustrated in Figs. 4(b), 4(c), and 4(d). Because the energy levels $E_{HO}$ and $E_{LU}$ exhibit very high probability densities at the zigzag edge sites, specifically at $j=1$ and $j=N_{a}$ (see Fig. 5), $V_{y}$-dependent $E_{HO}$ and $E_{LU}$ can be readily understood in terms of $\delta_{A}$ and $\delta_{B}$ as illustrated in Fig. 3. For example, $E_{LU}$ and $E_{HO}$ very close to CNP at $\delta=0.45$ eV and $V_{y}=0.9$ V, as shown in Fig. 4(b). The range of energy level degeneracy with respect to transverse electric fields aligns with the electronic structures shown in Fig. 3. Notably, while ZGNRs without staggered sublattice potential have demonstrated valley structures under strong transverse electric fields, the extent of level degeneracy is limited, and the electric field-direction-dependent valley structure disappears [LeeKW, 86].
@@
Since the transmission coefficient spectra of nanostructures are influenced by the wave-functions of their energy levels [Kuo3, 35], understanding the wave functions of each energy level for 4-AB/4-C/4-AB ZGNR segments with various $\delta$ values is crucial. In Fig. 5(a), the probability densities $|\Psi(\ell,j)(E=0)|^{2}$ demonstrate localized and symmetrical zigzag edge states in the $y$ direction for $\delta=0$ eV and $V_{y}=0$. Figures 5(b) and 5(c) show the probability densities of $E_{HO}=-0.253$ eV and $E_{LU}=0.253$ eV, respectively, depicting the wave functions of the HOMO and the LUMO at $\delta=0.27$ eV. Analysis of their probability density distributions reveals that $E_{HO(LU)}=\mp 0.253$ eV are primarily determined by $E_{B}$ and $E_{A}$, respectively. The maximum probability densities $|\Psi(\ell,j)(E_{HO})|^{2}$ and $|\Psi(\ell,j)(E_{LU})|^{2}$ are concentrated at $j=12$ and $j=1$, respectively, with $|\Psi(\ell,j)(E_{HO})|^{2}$ being vanishingly small for $j=7,9$ and $11$, and $|\Psi(\ell,j)(E_{LU})|^{2}$ for $j=2,4$ and $6$. In Fig. 5(d), 5(e) and 5(f) at $\delta=0.9$ eV, we present the probability density distributions of $E_{1}=-0.704$ eV, $E_{2}=-0.7053$ eV and $E_{3}=-0.7284$ eV, which are below the CNP. These states exhibit delocalized and asymmetrical characteristics in the $y$ direction, indicative of band inversion in flat bands for ZGNRs. The probability densities $|\Psi(\ell=1,j)(E_{i})|^{2}$ and $|\Psi(\ell=81,j)(E_{i})|^{2}$ significantly influence the transmission coefficient spectra of 4-AB/4-C/4-AB ZGNR segments with armchair edges coupled to the electrodes shown in Fig. 1(d).
@@
According to Eq. (2), the thermoelectric quantities of ZGNRs, including electrical conductance, Seebeck coefficient, and electron thermal conductance, are determined by the area and shape of the transmission coefficient (${\cal T}_{LR}(\varepsilon)$) curve. Therefore, it is necessary to clarify the effects of contact properties on ${\cal T}_{LR}(\varepsilon)$. In Fig. 6, we present the calculated ${\cal T}_{LR}(\varepsilon)$ curves of 4-AB/4-C/4-AB ZGNR segments for various $\Gamma_{t}$ values, which depend on the contact geometries and the electrode materials. For simplicity of analysis, $\varepsilon\leq 0$ is considered throughout this article. Due to electron-hole symmetry, the curves of transmission coefficients for $\varepsilon\geq 0$ are readily obtained (or see Fig. B.1). As seen in Fig. 6(a), the ${\cal T}_{LR}(\varepsilon)$ curve shows many discrete peaks under a weak coupling strength of $\Gamma_{t}=0.27$ eV. With increasing $\Gamma_{t}$ values, the area of ${\cal T}_{LR}(\varepsilon)$ curve increases, and the shapes of these areas also change. For instance, in Fig. 6(b), the area of ${\cal T}_{LR}(\varepsilon)$ curve shows a right-triangle shape with a steep change with respect to $\varepsilon$ on the side toward the central gap. In Figs. 6(c) and 6(d), the areas of ${\cal T}_{LR}(\varepsilon)$ curves show rectangular shapes. For $\Gamma_{t}=2.7$ eV, we observe an arch shape in Fig. 6(f). The level degeneracy is characterized by ${\cal T}_{LR}(\varepsilon)$ with values near two. For $\Gamma_{t}=2.7$ eV, the electrode materials can be regarded as graphenes since $\Gamma_{t}=t_{pp\pi}$. Because the spectra of the transmission coefficient ${\cal T}_{LR}(\varepsilon)$ involve the Green’s functions, the local density of states determined by the imaginary part of $G^{r(a)}(\varepsilon)$ directly reveals the $\Gamma_{t}$-dependent probability density of each energy level in Fig. 6, which determines the shape of the transmission coefficient curve. We refer readers to our previous work[Kuo3, 35].
@@
Figure 6: Transmission coefficient ${\cal T}_{LR}(\varepsilon)$ of 4-AB/4-C/4-AB ZGNR segments as functions of $\varepsilon$ for various $\Gamma_{t}$ values at $N_{a}=12$, $N_{z}=81$, $\delta=0.9$ eV and $V_{y}=-0.9$ V: (a) $\Gamma_{t}=0.27$ eV, (b) $\Gamma_{t}=0.45$ eV, (c) $\Gamma_{t}=0.63$ eV, (d) $\Gamma_{t}=0.81$ eV, (e) $\Gamma_{t}=0.99$ eV, and (f) $\Gamma_{t}=2.7$ eV. Here, $N_{z}=81$ corresponds to a channel length of $L_{z}=9.84$ nm.